OpenStudy (anonymous):

One canned juice drink is 15% orange juice; another is 5% orange juice. How many liters of each should be mixed together in order to get 10L that is 11% orange juice?

hero (hero):

Hey, I lost track of you.

OpenStudy (dumbcow):

It uses elimination again Let A be 15% juice and B is 5% juice A+B = 10 0.15A + 0.05B = 0.11*10 = 1.1 Multiply 2nd equation by 100 to get rid of decimals A+B = 10 15A + 5B = 110

OpenStudy (anonymous):

Okay, that went right over my head...LOL

hero (hero):

dumbunny, you're like me with Calculus

OpenStudy (anonymous):

haha I wouldn't touch Calculus with a 100 foot pole!

hero (hero):

Of course I aced Calculus I and II, classroom courses...still had issues though

hero (hero):

I would never have been able to pull off what some people do here.

OpenStudy (anonymous):

I am in awe of about 99 percent of the people here.

OpenStudy (anonymous):

So how many liters of the 15% orange juice should be in the mixture and how many liters of the 5% orange juice should be in the mixture?

OpenStudy (dumbcow):

ok sorry i was away A is amount of drink that is 15% orange juice B is amount of drink that is 5% orange juice So 0.15A and 0.05B represent the actual amount of orange juice in either drink. By mixing them together you get a total amount 10L You want this mixture to have 11% orange juice...so 0.11*10 = 1.1 Now set up balanced equations First equation is concerned with quantity of drink A + B = 10 2nd equation is concerned with actual amount of orange juice 0.15A + 0.05B = 1.1 Think of it this way, if there is more 15% drink then mixture will be closer to 15% than 5%, If there is more 5% used then mixture will be closer to 5% than 15% does that make more sense?

OpenStudy (anonymous):

No it doesn't. Sorry, I know I am pretty pathetic when it comes to this stuff.

OpenStudy (dumbcow):

ok lets do an example: Lets say we mix 100L of each drink. 15% of 100 = 0.15*100 = 15 5% of 100 = 0.05*100 = 5 This means we are mixing 15L of orange juice with 5L of orange juice total amount of mixture: 100 + 100 = 200 total amount of orange juice 15 + 5 = 20 Now to find out the percentage of orange juice in mixture, divide 20/200 = 1/10 = 0.1 = 10% Ok so if the question was how many liters to be mixed to get 200L of 10% orange juice. the answer would be 100 of A and 100 of B.

OpenStudy (anonymous):

So you would mix equal amounts? No, that can't be right. My mind is just a blank right now. I have been working on this quiz for the past 8 hours.

OpenStudy (dumbcow):

yes if you wanted mixture to be 10% which is exactly in between 5% and 15%. This problem is looking for 11% is 11 closer to 15 or 5?

OpenStudy (anonymous):

15

OpenStudy (dumbcow):

right, so you know that there will be more of the 15% in the mixture They have to add up to 10, half of 10 is 5 so one must be greater than 5 while the other less than 5

OpenStudy (anonymous):

6 and 4

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

So 6 liters of the 5% and 4 liters of the 15%

OpenStudy (dumbcow):

for this one it works out nicely, 6 and 4 most times you need to go through and solve the equations to get exact answer check your answer does 0.15*6 + 0.05*4 = 1.1 ??

OpenStudy (anonymous):

yes it does

OpenStudy (dumbcow):

then thats your solution :) wait which one should have more, 15% or 5% ?

OpenStudy (anonymous):

the 5%

OpenStudy (dumbcow):

is it closer to 11% ?

OpenStudy (anonymous):

no

OpenStudy (dumbcow):

whichever one is closer to the mixture percentage will always have more

OpenStudy (anonymous):

So it would be 6 liters of the 15% and 4 liters of the 5% ??

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

okay, got it. THANK GOD! Thank you very much.

OpenStudy (dumbcow):

your welcome